# CERN-TH-2024-177

### v1

@@ IV Bootstrap Fit results
using experimental and lattice phase shifts $\delta_{i}^{\text{exp}}$, where the index $i$ stands collectively for the quantum numbers $(I,\ell)=\{(0,0),(1,1),(2,0),(0,2)\}$, and the energy $s_{i}$ of the measurement. Our ansatz for the phase shift is obtained by projecting $\mathcal{A}^{\text{ansatz}}_{\Theta}$ in partial waves and using the definition $S_{\Theta}=\eta_{\Theta}\,\text{exp}(2i\delta_{\Theta})$.
@@
The result of the Bootstrap fit is shown in figure 3. The orange data points are experimental, the green are taken from the lattice. The dark blue curve is the best fit with minimum $\chi^{2}(\Theta)\approx 40$. We do not assign a statistical meaning to the value of the $\chi^{2}$, but take it as a likelihood measure. The largest contribution to the $\chi^{2}$ comes from the $P$ wave, with $\chi^{2}\approx 30$, where the experimental error is almost negligible. We plan to investigate this channel with more care in the future, replacing the phase shift data with the determination from the pion form factors de Troconiz and Yndurain (2005), and with lattice extrapolations Boyle et al. (2024). The second largest contribution of order ten is due to the $S_{0}$ wave. Here we observe a systematic error in reproducing the phase above the $K\bar{K}$ threshold $s\approx 50$ (the $\chi^{2}$ below this energy is of order one). We believe this is due to the lack of an explicit threshold in our ansatz at that scale, as corroborated by the slower convergence of the Bootstrap problem, see also Appendix A. Both $S_{2}$ and $D_{0}$ waves have $\chi^{2}$ of order one.
@@ V.1 Amplitude below threshold
Figure 4: Partial wave amplitudes $t_{\ell}^{(I)}$ as function of $s/m_{\pi}^{2}$ in the real strip $40<s<$ 4.
@@ V.2 The $D_{2}$ partial wave
Figure 5: Top figure, elasticity for the $D_{2}$ wave as function of $s/m_{\pi}^{2}$. The black line is the parametrization used in Pelaez and Yndurain (2005). Bottom figure, phase shift for the $D_{2}$ wave, and comparison with the available experimental data. Green and orange points are the two possible determinations for the phase taken from Hoogland et al. (1977).
@@ V.3 Spectrum
$m_{\sigma}$ (2.97–3.01)+$i$(2.07–2.11) (2.9–5.8)+$i$(0.7–5.8)
$m_{\rho^{\prime}}$ (8.9–10.1)+$i$(0.5–0.6) (10.4–11)+$i$(1.4–1.5)
$m_{\rho_{3}}$ (16.1–17.3)+$i$(0.51–0.85) (12.2–12.4)+$i$(0.63–0.73)
$m_{\mathbb{T}}$ (14.3–14.9)+$i$(2.3–2.5) ?
@@ V.4 The tetraquark
Looking at the $S_{2}$ wave, we find a surprise. If we continue to high energy, figure 9, we observe a broad resonance with isospin $I=2$. Such a state would be a genuine tetraquark. In the real world, if we reintroduce the electric charge, such particles should also carry a charge two. Its mass and width in physical units are around 2 GeV and 600 MeV. It might be interesting to look at the invariant mass distributions of the $\pi^{\pm}\pi^{\pm}$ pairs in decays that leave enough energy to reach the 2 GeV scale, e.g. $B^{+}\to(D^{-}/K^{-},/\pi^{-})\pi^{+}\pi^{+}$ or the charge conjugate ones.
Curiously, we know an example in two dimensions where this scenario is qualitatively realized. This is the two-dimensional theory of the QCD flux-tube. There, the role of the QCD $\rho$ is played by the axion, and the role of the $\sigma$ is played by the dilaton. In this model, the analogy is striking because we have a sharp axion and a broad dilaton Gaikwad et al. (2024). In Elias Miró et al. (2019a) it was noticed that crossing symmetry enforces a broad resonance in the symmetric channel. A similar mechanism might be responsible for the presence of the tetraquark observed in our amplitude.
This zero does not suffer from Bootstrap systematics. However, we have not imposed any inelasticity in this region, and inelastic effects might slightly modify the position of this resonance. It would be interesting to test this hypothesis and make our prediction stronger.
Figure 9: High energy $S_{2}$ partial wave phase shift profile. In blue the best fit, in light blue all amplitudes with suboptimal $\chi^{2}$. We see a jump in the phase around $s\approx 200$ that signals the presence of a resonance. In the inset, the parameters are extracted from the position of the zero.
@@ V.5 High Energy
Figure 10: Total cross sections for $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ and $\pi^{-}\pi^{-}\to\pi^{-}\pi^{-}$. In blue the best fit for $N_{\text{vars}}=547$, in light blue the best fit for $N_{\text{vars}}=397$. Above 2 GeV there is still room to improve numerics, but there is already a very good agreement between Bootstrap and the experimental data Biswas et al. (1967); Robertson et al. (1973); Cohen et al. (1973); Losty et al. (1974); Hanlon et al. (1976); Abramowicz et al. (1980); Zakharov and Sergeev (1984).
Finally, we discuss the Regge behavior of our amplitude and compare it with experimental data.
The emergent high-energy behavior of S-matrix Bootstrap amplitudes is the most surprising aspect recently understood. Although the Bootstrap ansatz in (3) goes to a constant at infinity, there is an intermediate regime in which the numerical amplitudes behave according to Regge theory, showing at large $s$ and fixed $t$ the expected behavior $\mathcal{A}\sim s^{\alpha(t)}$ Bhat et al. (2024); Gumus (2023); Elias Miró et al. ; Correia et al. .
In figure 10, we compare the experimental total cross sections for $\pi^{+}\pi^{-}$ and $\pi^{-}\pi^{-}$ with those extracted from our best-fit amplitude in blue [^21] In isospin components $\sigma_{\pi^{+}\pi^{-}}\propto\tfrac{1}{6}(2\mathcal{T}^{(0)}+3\mathcal{T}^{(1)}+\mathcal{T}^{(2)})$, and $\sigma_{\pi^{-}\pi^{-}}\propto\mathcal{T}^{(2)}$.
@@
The solid blue curve is from our fit with the highest $N_{\text{vars}}=547$, while the light blue is the best fit with $N_{\text{vars}}=397$.
In the $\sigma_{\pi^{+}\pi^{-}}$ we can observe nice peak structures corresponding to the various resonances in the spectrum (the first and highest is the $\rho$ peak). In the $\sigma_{\pi^{-}\pi^{-}}$ we do not expect any peak due to resonances, except for the tetraquark around 2 GeV. At asymptotically high energies, we expect a constant behavior compatible with the pomeron exchange. As we increase $N_{\text{vars}}$ we expect the Bootstrap amplitude to better approximate this behavior which is not automatically imposed in our ansatz (3).
@@ VI Discussion and outlook
This Letter proposes a novel constrained approach to fit experimental data. From our perspective, the advantage we foresee compared to other methods is the control of the theoretical approximations. The amplitude model used to fit the data is a genuine analytic and crossing symmetric function compatible with non-perturbative unitarity. The algorithm mixes semi-definite and non-convex optimization techniques: for the Bootstrap part we use SDPB Simmons-Duffin (2015); Landry and Simmons-Duffin (2019), and for the PSO we have coded a simple Mathematica notebook.
The $\pi\pi\to\pi\pi$ amplitude constructed has many interesting features. It reproduces the experimental data, correctly predicts the whole $G$-parity plus spectrum below 1.4 GeV, agrees with the $\chi$PT at low energy, and with the high-energy regime. We correctly predict an additional resonance heavier than the $\rho$ in the $P$ wave with a comparable mass to the real world $\rho(1450)$, and a spin three resonance 20% heavier than the experimental result. The $D_{2}$ phase shift and inelasticity agree with experimental data and phenomenological models.
@@
Surprisingly, we predict a tetraquark state with a mass of around 2 GeV in the $I=2$ channel. This is a nonperturbative prediction, and we are not aware of any other model making the same statement. Given that the amplitude constructed correctly describes and predicts several key features of $\pi\pi$ scattering amplitudes, we are quite confident this particle should be there, and it would be interesting to look for it in the decays of $B$ mesons.
There are many improvements we can implement to obtain a better model for the pion amplitude.
@@
First, it would be interesting to include the correlation matrix for the lattice data Blum et al. (2021) in the $S_{0}$ wave, and investigate their effect on our determination of the $f_{0}(500)$ mass parameters. For the $P$ we could also include the $\delta_{1}^{(1)}$ phase extracted from form factors de Troconiz and Yndurain (2005) and the lattice QCD extrapolation at the physical pion mass Boyle et al. (2024).
The kink functional (10) used to construct our fit model may seem ad hoc and not easy to generalize to other physical systems. For this reason, we plan to repeat the fit using different optimization problems Guerrieri et al. . The existence of alternative constructions can also be considered a source of systematic that should be included in our estimates, which we currently do not account for. [^22] We have already tested a different fit approach which gives almost identical results.
We believe replacing the spectrum assumption and generating it dynamically would be ideal. This could be done by optimizing larger linear combinations of low-energy constants, generalizing the kink functional in equation (10), as it was already done in several S-matrix Bootstrap works Elias Miró et al. (2019b); Guerrieri et al. (2021a); Bose et al. (2021); Guerrieri et al. (2021b, 2023b); Häring et al. (2024); Acanfora et al. (2024).
It would also be interesting to analytically continue the best-fit solution in spin, as done in Acanfora et al. (2024), and examine the complexified Regge trajectories. This would require running with a larger ansatz to have enough numerical precision. We could investigate whether the three $0^{+}(0^{++})$ scalar resonances below 1.4 GeV are subtraction poles or part of Regge trajectories.
@@
Another systematic of our fit procedure is the input of the shape of the inelasticity. To make the fit self-consistent we could enlarge the set of amplitudes including mixed processes involving pions and kaons. By doing so, we would have access to many additional resonances, and we could use more experimental data as input. The inelasticity profile would be an output of the fit.
@@
It would also be interesting to use our amplitude to measure form factors, see Karateev et al. (2020); Correia et al. (2023); Cordova et al. (2024). Fixing our parametrization, we could read off the form factors as an output, and check consistency with high-energy sum rules recently discussed in He and Kruczenski (2023, 2024) for the case of $\pi\pi$ scattering. [^23] Knowledge of the form factors computed using other methods can also be used to constrain the amplitude, see e.g. Chen et al. (2022)., where the authors used Hamiltonian truncation to compute them.
@@
Ultimately, how can we extend our method and determine rigorous bounds on the position of the resonances in figure 1? It would be interesting in this sense to generalize the strategy proposed in Gabai and Yin (2022), where the authors obtained bounds on the position of resonances in the complex $s$ plane in the case of Ising Field Theory in $1+1$ dimensions.
@@
Finally, we could explore processes involving other particles rather than pions. For example, Glueball couplings determine the dependence of the lattice energy levels on the Glueball-Glueball phase shift in a non-perturbative way. Combining the scattering amplitudes constructed in Guerrieri et al. (2023a) with the lattice energy levels we could learn about observables that may seem hopeless to measure.
@@ Acknowledgements
We thank Mattia Bruno, Gilberto Colangelo, Miguel Correia, Barak Gabai, Victor Gorbenko, David Gross, Aditya Hebbar, Denis Karateev, Harish Murali, Jose Ramon Pelaez, Joao Penedones, Alessandro Pilloni, Jiaxin Qiao, Riccardo Rattazzi, Slava Rychkov, Balt van Rees, David Simmons-Duffin, Pedro Vieira, and Alexander Zhiboedov for useful discussions.
We thank Joao Penedones, Alessandro Pilloni, Balt van Rees, Pedro Vieira, Mark Wise and Alexander Zhiboedov for the many useful comments on the draft.
@@
Figure 12: Real (blue), imaginary (red), and absolute value (green) of the phase shifts for various channels. Thick lines are from our best numerics with $(N,M)=(14,12)$, light lines for $(N,M)=(12,10)$. The inelasticity profiles used as a constraint are denoted in black. We do not draw it for the channels where we impose the simple probability conservation $|S_{J}^{(I)}|\leq 1$.

### v2

@@ IV Bootstrap Fit results
using experimental and lattice phase shifts $\delta_{i}^{\text{exp}}$, where the index $i$ stands collectively for the quantum numbers $(I,\ell)=\{(0,0),(1,1),(2,0),(0,2)\}$, and the energy $s_{i}$ of the measurement. Our ansatz for the phase shift is obtained by projecting $\mathcal{A}^{\text{ansatz}}_{\Theta}$ in partial waves and using the definition $S_{\Theta}=|S_{\Theta}|\,\text{exp}(2i\delta_{\Theta})$.
@@
The result of the Bootstrap fit is shown in figure 4. The orange data points are experimental, the green are taken from the lattice. The dark blue curve is the best fit with minimum $\chi^{2}(\Theta)\approx 40$. We do not assign a statistical meaning to the value of the $\chi^{2}$, but take it as a likelihood measure. The largest contribution to the $\chi^{2}$ comes from the $P$ wave, with $\chi^{2}\approx 30$, where the experimental error is almost negligible. We plan to investigate this channel with more care in the future, replacing the phase shift data with the determination from the pion form factors de Troconiz and Yndurain (2005); Colangelo et al. (2019), and with lattice extrapolations Boyle et al. (2024). The second largest contribution of order ten is due to the $S_{0}$ wave. Here we observe a systematic error in reproducing the phase above the $K\bar{K}$ threshold $s\approx 50$ (the $\chi^{2}$ below this energy is of order one). We believe this is due to the lack of an explicit threshold in our ansatz at that scale, as corroborated by the slower convergence of the Bootstrap problem, see also Appendix A. Both $S_{2}$ and $D_{0}$ waves have $\chi^{2}$ of order one.
@@ V.1 Amplitude below threshold
Figure 5: Partial wave amplitudes $t_{\ell}^{(I)}$ in the real strip $40<s<$ 4. The black line is the Bootstrap prediction, dashed lines are different approximations from $\chi$PT Bijnens et al. (1996).
@@ V.2 The $D_{2}$ partial wave
Figure 6: Top panel: Elasticity of the $D_{2}$ wave as a function of $s/m_{\pi}^{2}$. The black line represents the parametrization from Pelaez and Yndurain (2005). Bottom panel: Phase shift of the $D_{2}$ wave compared with the available experimental data. The green and orange points correspond to the two possible phase determinations reported in Hoogland et al. (1977).
@@ V.3 Spectrum
$m_{\sigma}$ $(2.99{\pm}0.02)$+$i(2.09{\pm}0.02)$ $(4.3\pm 1.5)$+$i(3.3\pm 2.6)$
$m_{\rho^{\prime}}$ $(9.5\pm 0.6)$+$i(0.55\pm 0.05)$ $(10.7\pm 0.3)$+$i(1.45\pm 0.05)$
$m_{\rho_{3}}$ $(16.7\pm 0.6)$+$i(0.7\pm 0.2)$ $(12.3\pm 0.1)$+$i(0.68\pm 0.05)$
$m_{\mathbb{T}}$ $(14.6\pm 0.3)$+$i(2.4\pm 0.1)$
@@ V.4 The tetraquark
Figure 10: Real and imaginary parts of the amplitude in the $S_{0}^{(2)}$ channel. The error bars indicate the systematic uncertainties in the line shape of this function arising from the unknown inelasticity above 1.4 GeV. This error is comparable to the statistical error of our fit reported in figure 2.
As already indicated in the introduction, examining the $S_{2}$ wave reveals an unexpected result. Continuing the phase shift $\delta_{0}^{(2)}$ to high energy, as shown in figure 2, we observe a broad resonance with isospin $I=2$. This indicates the presence of a genuine Tetraquark state Jaffe (1977); ’t Hooft et al. (2008). In the physical world, where electric charge is reintroduced, such particles would carry a charge of two. Its mass and width in physical units are respectively 2 GeV and 600 MeV. It would be intriguing to examine the invariant mass distributions of $\pi^{\pm}\pi^{\pm}$ in $B$-meson decays that provide sufficient energy to reach the 2 GeV scale. Examples include the decay $B^{+}\to\pi^{-}\pi^{+}\pi^{+}$ Aaij et al. (2020), which can be studied at LHCb, as well as similar decays where the $\pi^{-}$ is replaced by either $D^{-}$ or $K^{-}$.
We determine the mass and width of the Tetraquark by locating the zero in the complex $s$ plane in the $S_{2}$ wave. This determination is free from Bootstrap-related systematics, as the amplitude in this channel is well-converged. However, we have not accounted for inelastic effects in this region, which could slightly alter the resonance’s position.
In figure 10, we plot the real and imaginary parts of the partial wave amplitude $t_{0}^{(2)}$ as a function of $\sqrt{s}$ measured in GeV. From an amplitude perspective, this represents the experimental signal expected in the presence of a Tetraquark. [^21] We thank Alessandro Pilloni for providing insights into the input required for the experimental data analysis. The black dashed line separates the region where we fit the data and incorporates the inelasticity profile from the high-energy region, which is unconstrained. The gray curve represents the elasticity profile $\eta_{0}^{(2)}(s)$ taken from Garcia-Martin et al. (2011). Above $\sqrt{s}=1.4$ GeV, we set $\eta_{0}^{(2)}(s)=1$ in our procedure. For illustrative purposes, we show the amplitude profile using the $S_{2}$ wave parametrization $S_{0}^{(2)}(s)=|S_{0}^{(2)}(s)|\exp(2i\delta_{0}^{(2)}(s))$, where $\delta_{0}^{(2)}$ is the phase obtained from our best fit, and $|S_{0}^{(2)}(s)|$ is a generic function bounded by $\eta_{0}^{(2)}(s)\leq|S_{0}^{(2)}(s)|\leq 1$, for $s>1.4$ GeV. The lower bound corresponds to the extrapolation of Garcia-Martin et al. (2011) to higher energies, as shown by the gray area in figure 10. The blue and red bands depict how uncertainty in the elasticity function influences the shape of the amplitude.
It would be interesting to understand the mechanism behind the emergence of the Tetraquark, using, for instance, the Roy equations Roy (1971). Notably, there is a two-dimensional example that qualitatively mirrors this scenario: the two-dimensional theory of the QCD flux tube. In this context, the QCD $\rho$ is analogous to the axion, while the $\sigma$ corresponds to the dilaton. The analogy is striking, as this model features a sharp axion and a broad dilaton Gaikwad et al. (2024). Furthermore, in Elias Miró et al. (2019a), it was observed that crossing symmetry necessitates a broad resonance in the symmetric channel. A similar mechanism might underlie the presence of the Tetraquark identified in our amplitude.
@@ V.5 High Energy
Figure 11: Total cross sections for $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ and $\pi^{-}\pi^{-}\to\pi^{-}\pi^{-}$. The gray shaded area is the region where we fit the data. The blue curve is the best fit for $N_{\text{vars}}=547$, in light blue the best fit for $N_{\text{vars}}=397$. Above 2 GeV there is still room to improve numerics, but there is already a very good agreement between Bootstrap and the experimental data Biswas et al. (1967); Robertson et al. (1973); Cohen et al. (1973); Losty et al. (1974); Hanlon et al. (1976); Abramowicz et al. (1980); Zakharov and Sergeev (1984).
Finally, we examine the Regge behavior of our amplitude and compare it with experimental data.
One of the most intriguing recent insights into S-matrix Bootstrap amplitudes is their emergent high-energy behavior. Although the Bootstrap ansatz in (3) approaches a constant at infinity, there exists an intermediate regime where the numerical amplitudes align with Regge theory. Specifically, at large $s$ and fixed $t$, the amplitudes exhibit the expected behavior $\mathcal{A}\sim s^{\alpha(t)}$ Bhat et al. (2024); Gumus (2023); Elias Miró et al. ; Correia et al. .
In figure 11, we compare the experimental total cross sections for $\pi^{+}\pi^{-}$ and $\pi^{-}\pi^{-}$ with those extracted from our best-fit amplitude, shown in blue [^22] In isospin components $\sigma_{\pi^{+}\pi^{-}}\propto\tfrac{1}{6}(2\mathcal{T}^{(0)}+3\mathcal{T}^{(1)}+\mathcal{T}^{(2)})$, and $\sigma_{\pi^{-}\pi^{-}}\propto\mathcal{T}^{(2)}$.
@@
The solid blue curve represents our fit with the highest $N_{\text{vars}}=547$, while the light blue curve corresponds to the best fit with $N_{\text{vars}}=397$.
In $\sigma_{\pi^{+}\pi^{-}}$ we observe distinct peak structures corresponding to various resonances in the spectrum, with the first and most prominent being the $\rho$ peak. In contrast, $\sigma_{\pi^{-}\pi^{-}}$ is not expected to exhibit resonance-related peaks, except for the Tetraquark. At asymptotically high energies, we expect a constant or slightly growing behavior consistent with the pomeron exchange (see for instance Caprini et al. (2012)). As $N_{\text{vars}}$ increases, we expect the Bootstrap amplitude will better approximate this behavior, which is not explicitly enforced in our ansatz (3).
@@ VI Discussion and outlook
This Letter introduces a novel constrained approach for fitting experimental data. The key advantage of our method lies in its control over theoretical approximations. The amplitude model employed is a genuinely analytic, crossing-symmetric function that satisfies non-perturbative unitarity. The algorithm combines semi-definite programming and non-convex optimization techniques: for the Bootstrap component, we use SDPB Simmons-Duffin (2015); Landry and Simmons-Duffin (2019), while for the particle swarm optimization (PSO), we implemented a straightforward Mathematica notebook.
The constructed $\pi\pi\to\pi\pi$ amplitude exhibits several remarkable features. It reproduces the experimental data, accurately predicts the $G$-parity plus spectrum below 1.4 GeV, aligns with $\chi$PT at low energies, and qualitatively matches the expected behavior in the high-energy regime. Additionally, the model predicts a resonance heavier than the $\rho$ in the $P$ wave, with a mass comparable to the real-world $\rho(1450)$, as well as a spin-three resonance approximately 20% heavier than the experimental value. The $D_{2}$ phase shift and inelasticity also show good agreement with experimental data and phenomenological models.
@@
Interestingly, our analysis predicts a tetraquark state with a mass of approximately 2 GeV in the $I=2$ channel. This is a nonperturbative prediction, and to our knowledge, no other model makes a similar claim. Given that the constructed amplitude accurately describes and predicts several key features of $\pi\pi$ scattering, we are confident in the existence of this particle. Investigating its presence in the decays of $B$ mesons would be particularly intriguing.
Several enhancements could further improve the pion amplitude model we have constructed.
@@
First, incorporating the correlation matrix for the lattice data in the $S_{0}$ wave from Blum et al. (2021) would be valuable for assessing its impact on the determination of the $f_{0}(500)$ mass parameters. In particular, the lattice data are the source of the small tension with the previous $\sigma$ determination from Roy equations Caprini et al. (2006). For the $P$ wave, it would be worthwhile to include the $\delta_{1}^{(1)}$ phase extracted from form factors de Troconiz and Yndurain (2005); Colangelo et al. (2019) as well as results from lattice QCD extrapolations at the physical pion mass Boyle et al. (2024).
The functional (10) used to construct our fit model might appear ad hoc and challenging to generalize to other physical systems. To address this, we plan to explore alternative optimization approaches in future work Guerrieri et al. . The availability of such alternative constructions could also introduce a source of systematic uncertainty that should be incorporated into our estimates—something not currently accounted for. [^23] We have already tested an alternative fitting approach, which yields nearly identical results. It would also be interesting to understand the relation between this objective and the regions selected in Bose et al. (2020, 2021), and He and Kruczenski (2023, 2024) using alternative principles.
@@
Another source of systematic uncertainty in our fit procedure is the input for the shape of the inelasticity. To make the fit more self-consistent, we could expand the set of amplitudes to include mixed processes involving pions and kaons. This would provide access to additional resonances and allow us to incorporate more experimental data. The inelasticity profile could then be treated as an output of the fit.
Moreover, we believe that replacing the spectrum assumption and dynamically generating all resonances would be ideal. This could be achieved by optimizing larger linear combinations of low-energy constants and generalizing the functional in equation (10) as has been done in several S-matrix Bootstrap studies Elias Miró et al. (2019b); Guerrieri et al. (2021a); Bose et al. (2021); Guerrieri et al. (2021b, 2023b); Häring et al. (2024); Acanfora et al. (2024).
It would also be interesting to analytically continue the best-fit amplitude in spin, as done in Acanfora et al. (2024), and explore the Regge trajectories for complex spins. This would necessitate using a larger ansatz to ensure sufficient numerical precision. We could then investigate in detail the resonance spectrum, and understand whether the three $0^{+}(0^{++})$ scalar resonances below 1.4 GeV are subtraction poles or part of Regge trajectories.
@@
Our amplitude could also be used to measure form factors, as explored in Karateev et al. (2020); Correia et al. (2023); Cordova et al. (2024). By fixing our parametrization, we could extract the form factors as an output and verify their consistency with high-energy sum rules recently explored in He and Kruczenski (2023, 2024) for $\pi\pi$ scattering. [^24] Form factors computed using other methods, such as in Chen et al. (2022), where Hamiltonian truncation was employed, could also be used to constrain the amplitude.
@@
Ultimately, how can we extend our method and determine rigorous bounds on the position of the resonances in figure 1 determined from the Bootstrap Fit procedure? In this direction, it would be interesting to generalize the strategy proposed in Gabai and Yin (2022), where the authors obtained bounds on the position of resonances in the complex $s$ plane in the case of Ising Field Theory in $1+1$ dimensions.
@@
Finally, we could explore processes involving particles other than pions. For instance, Glueball couplings influence the lattice energy levels through the Glueball-Glueball phase shift in a non-perturbative manner. By combining the scattering amplitudes constructed in Guerrieri et al. (2023a) with lattice energy levels, we could gain insights into observables that may otherwise be difficult to measure using standard tools.
@@ Acknowledgements
We thank Mattia Bruno, Gilberto Colangelo, Miguel Correia, Barak Gabai, Victor Gorbenko, David Gross, Aditya Hebbar, Denis Karateev, Luciano Maiani, Harish Murali, Jose Ramon Pelaez, Joao Penedones, Alessandro Pilloni, Jiaxin Qiao, Riccardo Rattazzi, Slava Rychkov, Balt van Rees, David Simmons-Duffin, Pedro Vieira, and Alexander Zhiboedov for useful discussions.
We thank Gilberto Colangelo, Joao Penedones, Alessandro Pilloni, Balt van Rees, Pedro Vieira, Mark Wise and Alexander Zhiboedov for the many useful comments on the draft.
@@
Figure 13: Real (blue), imaginary (red), and absolute value (green) of the phase shifts for various channels. Thick lines are from our best numerics with $(N,M)=(14,12)$, light lines for $(N,M)=(12,10)$. The elasticity profiles used as a constraint are denoted in black. We do not draw it for the channels where we impose the simple probability conservation $|S_{J}^{(I)}|\leq 1$.